OpenStudy (kaleidoscopicsink):

Can someone explain to me how to implicitly derive x^2-y^2= xy? I got to 2x - 2y dy/dx = y+x.

Parth (parthkohli):

why the y + x?

OpenStudy (kaleidoscopicsink):

Product rule gives you that when deriving the right side of the equation.

Parth (parthkohli):

it doesn't say that the derivative of xy is x + y, does it?

OpenStudy (kaleidoscopicsink):

IT does though cause you do, CB + AD(I work in terms of A,B,C, and D) SO A is x, B is y C is the derivative of x so 1, and D is the derivative of y so 1. and when done in CB + AD you get y+x.

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} uv = uv' + vu'\)

OpenStudy (kaleidoscopicsink):

Yes, which is the same as my CB+AD.

OpenStudy (mathstudent55):

The derivative of a product is the first times the derivative of the second plus the second times the derivative of the first.

OpenStudy (mathstudent55):

The derivative of x with respect to x is 1. The derivative of y with respect to x is y'.

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} (xy) = xy' + yx' = xy' + y\)

OpenStudy (anonymous):

Hi, maybe you don't know how to derive a phrase containing both x and y. here's an example: \[x ^{2}y \rightarrow \frac{ d(x ^{2}y) }{ dx }={ 2xy+x^2y' }\]

OpenStudy (kaleidoscopicsink):

I know how to derive with x and y. I just keep making such simple and stupid mistakes.

OpenStudy (anonymous):

Ok, then. go ahead then and put it here to be checked if you want.

OpenStudy (mathstudent55):

Remember the power rule and the chain rule.

OpenStudy (mathstudent55):

\(x^2-y^2= xy\) \(2x - 2yy' = xy' + yx'\) Do you understand it up to here?

OpenStudy (kaleidoscopicsink):

Yes, of course, I'm just struggling to do the y' part that you posted. I think the answer is (y-2x)/ (-2y-x). And what mathstudent55 wrote about the product rule I have written done I just didn't post the y' part. My equation looks like this in my notebook: 2x-2y(dy/dx)= y+x*(dy/dx)

OpenStudy (mathstudent55):

Ok, here is the rest. \(xy' + 2yy' = 2x - y\) \(y'(x + 2y) = 2x - y\) \(y' = \dfrac{2x - y}{x + 2y} \)

OpenStudy (kaleidoscopicsink):

2x − 2y 2x/1+ 2y y - 2x/ -2y - x 0 Those are my options. I originally got what you are showing, but I thought it was wrong.

OpenStudy (mathstudent55):

\(y' = \dfrac{2x - y}{x + 2y}\) \(y' = \dfrac{-1(2x - y)}{-1(x + 2y)}\) \(y' = \dfrac{-2x + y}{-x - 2y}\) \(y' = \dfrac{y-2x}{-2y -x }\)

OpenStudy (kaleidoscopicsink):

Wow, so I unnecessarily confused myself, but had the right answer. Thank you everyone for all your help. cx And a huge thank you to @mathstudent55

OpenStudy (mathstudent55):

You are very welcome. You are correct. You did have the correct answer. You just had to make it look like one of the choices, so you could choose from the choices.